# Data-driven Energy Efficiency Modelling in Large-scale Networks: An Expert Knowledge and ML-based Approach

### v1

@@ VII-C3 User transfer
Transfer probability, $p^{\rm HO}_{c^{\rm a},j,h^{\prime}}$: Let us denote by $p^{\rm HO}_{c^{\rm a},j,h^{\prime}}$ the probability of a UE transferring from capacity cell, $c^{\rm a}$, to a neighboring cell, $j\in{\cal I}_{c^{\rm a}}$, when performing an inter-frequency handover through the A4 handover entry condition, at hour, hsuperscriptnormal-$h^{\prime}$.
@@ VII-D Rolling from hour to hour
Algorithm 1 summarised this fused Monte Carlo and ABM framework.
@@ VIII Experiments and Analysis
To evaluate the framework introduced in Section VII, we employ datasets from a real network in a metropolitan area of China. Our focus is on a set denoted as $\mathcal{C}$, comprising 657 cells categorized into (i) 375 capacity booster cells (constituting the set ${\cal C_{\rm C}}$) and (ii) 282 coverage cells (comprising the set ${\cal C_{\rm B}}$). These datasets spanned 12 days in April 2022 and were crucial for constructing the models detailed in Sections V, VI, VII, and VIII-A. Additionally, data from a 3-day period in May 2023 were utilized to create the test set and calculate the ground truth, vital for validating our model results.
It’s important to note that each dataset has been aggregated on an hourly basis, providing 24 values per day for each cell. These values offer insights into traffic, energy consumption, and measurement report statistics (see Table II and III).
@@ VIII-A Benchmark algorithm
VIII-B Modelling the shutting down time of capacity cells
@@
VIII-C Estimating the load of network cells
@@
VIII-D Modelling the energy consumption of capacity cells
@@
VIII-E Characterizing the UE DL Rate
@@ IX Conclusions
In conclusion, this paper introduced and evaluated the SRCON framework, a comprehensive data-driven approach for modeling the energy efficiency of cellular networks while consdering carrier shutdown policies. Leveraging advanced ML- and expert-based models, SRCON accurately estimated key network performance parameters, including carrier shutdown duration, cell load, energy consumption, and UE data rates. Notably, SRCON demonstrated remarkable accuracy in predicting these indicators in response to various network-wide carrier shutdown configurations.
The proposed ML-based energy consumption model, utilizing estimated carrier cell shutdown times and cell loads from our novel, expert-based ABM, achieved a substantial accuracy gain compared to a state-of-the-art energy consumption modeling approach used by a network operator, with a 62.08 % reduction in MAE and a 62.06 % reduction in MAPE. This underscores the effectiveness of our ABM in mimicking the underlying stochastic processes in the network.
Furthermore, the analysis of UE data rates showcased SRCON’s ability to capture the dynamic trade-offs between cell load, energy-saving policies, and UE rates. Compared to the benchmark data-driven estimator, SRCON achieved significant improvements, with approximately a 50 % reduction in MAE and a 57 % reduction in MAPE when estimating UE data rates.
In summary, SRCON represents a paradigm shift in practical network modeling and optimization, utilizing readily available network data to predict network resource use, energy consumption, and UE data rates. The results demonstrate the feasibility of modeling and optimizing wide-area networks without the need for expensive DT and/or ray tracing-based simulation tools.
@@ acknowledgment
generic access network

### v2

@@ VII-C3 User transfer
Transfer probability, $p^{\rm HO}_{c^{\rm a},j,h^{\prime}}$: Let us denote by $p^{\rm HO}_{c^{\rm a},j,h^{\prime}}$ the probability of a UE transferring from capacity cell, $c^{\rm a}$, to a neighboring cell, $j\in{\cal I}_{c^{\rm a}}$, when performing an inter-frequency handover through the A4 handover entry condition, at hour, $h^{\prime}$.
@@ VII-D Rolling from hour to hour
Algorithm 1 summarises this fused Monte Carlo and ABM framework.
@@ VIII Experiments and Analysis
To evaluate the framework introduced in Section VII, we employ datasets from a real network in a metropolitan area of China. Our focus is on a set denoted as $\mathcal{C}$, comprising 657 cells categorized into (i) 375 capacity booster cells (constituting the set ${\cal C_{\rm C}}$) and (ii) 282 coverage cells (comprising the set ${\cal C_{\rm B}}$). These datasets spanned 12 days in April 2022 and were crucial for constructing the models detailed in Sections V, VI, VII, and VIII-A. Additionally, data from a 3-day period in May 2023 were utilized to create the test set and calculate the ’ground truth’, vital for validating our model results. Ground truth denotes the actual measured truth, as measured directly from our dataset. It is important to note that each dataset, and thus our ground truth, has been aggregated on an hourly basis, providing 24 values per day for each cell. These values offer insights into traffic, energy consumption, and measurement report statistics (see Table II and III).
@@ VIII-A Benchmark algorithm
VIII-B Convergence analysis of ABM estimation errors
To validate the capabilities of the proposed model, this section analyzes the convergence of estimation errors as we increase the maximum number of time steps allowed in each Monte Carlo run of the ABM. Fig. 9 demonstrates that the average MAE of the estimates produced by the ABM is significantly influenced by this parameter. Notably, when this threshold is extended beyond 150, both types of errors rapidly decrease, achieving convergence once the maximum number of steps per Monte Carlo run reaches 400. By setting the maximum number of steps per Monte Carlo run to this level, ABM requires approximately 2 minutes to simulate network performance over a 24-hour period. This outcome lays the groundwork for discussing SRCON modelling performance in subsequent sections.
Figure 9: Average MAEs with respect to the maximum number of time steps per ABM iteration.
VIII-C Predicting the shutting down time of capacity cells
@@
VIII-D Predicting the load of network cells
@@
VIII-E Predicting the energy consumption of capacity cells
@@
VIII-F Predicting the UE DL Rate
@@
VIII-G Discussion on generalization to other scenarios
It is imperative to note that the precise calibration of our model is fundamentally intertwined with the specificities of the vendor’s equipment and the deployment characteristics it was originally designed for. Consequently, the model presented in this paper exhibits an optimized performance for metropolitan networks that deploy analogous technological frameworks than those used in training scenario, specifically 4G and 5G networks.
The necessity for comprehensive retraining of the ML models arises when attempting to transpose our energy-saving solutions and performance prediction models to alternative network setups or deployment scenarios. This retraining process is vital for accommodating the varying power consumption profiles and environmental conditions unique to different equipment or scenarios.
Furthermore, deviations in the carrier shutdown and handover (HO) algorithms, from the framework specified, necessitate a thoughtful redesign of the corresponding expert-based models within our ABM. Such adaptations underscore our commitment to providing precise, localized predictions.
This deliberation accentuates the importance of a thorough evaluation of network infrastructure, vendor-specific technologies, and the nuanced characteristics of the network environment prior to the deployment of our proposed model in real-world settings.
@@ IX Conclusions
In conclusion, this paper not only introduced the SRCON framework, but also rigorously evaluated its capabilities in the context of the research questions and scientific objectives laid out at the outset. Through this comprehensive data-driven approach, aimed at modeling the energy efficiency of cellular networks with a focus on carrier shutdown policies, we have made significant strides towards addressing the pressing challenges identified in our research.
Feasibility of Data-Driven Modeling: Our findings affirmatively answer the first research question, demonstrating that a data-driven model, relying exclusively on readily accessible network data and excluding costly DT, can indeed deliver accurate predictions of network energy consumption and UE throughput. The SRCON framework, by leveraging advanced ML- and expert-based models, has accurately estimated key network performance parameters in response to diverse carrier shutdown configurations, thereby showcasing remarkable accuracy.
Blending Machine Learning with Domain Expertise: In addressing our second research question, the SRCON framework exemplifies the seamless integration of ML algorithms with domain expertise. Our innovative, expert-driven ABM incorporates expert knowledge by implementing product algorithms, specifically focusing on carrier shutdown and HO mechanisms. It also effectively interacts with our novel ML-based models for energy consumption and UE throughput to ensure precise predictions. This integration has resulted in significant accuracy improvements, underscoring the efficacy of our approach in advancing energy efficiency modeling within large-scale networks.
Advancements and Comparative Evaluation: Regarding our third question, the SRCON framework has introduced notable advancements in network energy efficiency prediction. When compared to existing methodologies, notably a state-of-the-art approach used by a network operator, SRCON achieved a 62.08 % reduction in MAE and a 62.06 % reduction in MAPE for energy consumption modeling. For UE throughput estimation, significant improvements were observed with approximately a 50 % reduction in MAE and a 57 % reduction in MAPE, underscoring its superior predictive performance.
In summary, the SRCON framework signifies a paradigm shift towards practical, data-driven network modeling and optimization. It leverages existing network data to accurately predict resource utilization, energy consumption, and UE throughput, offering a viable solution for modeling and optimizing wide-area networks and fulfilling both the research questions posed and the scientific objectives set forth at the beginning of this study.
X Acknowledgement
This research is supported in part by the NSF project CNS 1910153, the Generalitat Valenciana through the CIDEGENT PlaGenT, Grant CIDEXG/2022/17, Project iTENTE, and by the action CNS2023-144333, financed by MCIN/AEI/10.13039/501100011033 and the European Union “NextGenerationEU”/PRTR.”
@@ acknowledgment
generative adversarial network